OpenStudy (kydawg):

Find all solutions to the equation. 7sin^2 x - 14sin x + 2 = -5

OpenStudy (mathstudent55):

Move the -5 to the left. Then you have a quadratic in sin x

OpenStudy (kydawg):

So you get 7sin^2 x -14sin x + 7 = 0 and that turns into x(7sin - 7)(sin - 1)?

OpenStudy (mathstudent55):

No. Treat sin x as if it were a variable. One method to avoid confusion is to do a substitution. Let \(u = \sin x\). Then \(u^2 = \sin^2 x\). Now rewrite your quadratic equation substituting the u's above for the sines.

OpenStudy (kydawg):

7u^2x - 14ux + 7 = 0 could you divide by 7 here?

OpenStudy (mathstudent55):

\(7\color{red}{\sin^2 x} - 14\color{green}{\sin x} + 7 = 0\) Substitutions: \(\color{red}{u^2 \sin^2 x }\) \(\color{green}{u = \sin x}\) You now have: \(7u^2 - 14u + 7 = 0\) and you have a simple quadratic equation in u. Now you solve. You can try factoring or the quadratic formula to solve for u.

OpenStudy (mathstudent55):

You are close, but notice that u = sin x, so when you substitute, there are no x's left. The u's take over completely as the variable.

OpenStudy (mathstudent55):

You are correct about dividing by 7. Take out the common factor of 7.

OpenStudy (kydawg):

(7u - 7)(u - 1) so u=1 you also get this if you divide by 7: (u - 1)(u - 1) u=1 in each

OpenStudy (mathstudent55):

Above in red it should read: \(\color{red}{u^2 = \sin^2 x}\) I accidentally left out the equal sign.

OpenStudy (mathstudent55):

Correct. The substitution allowed you to solve the quadratic and get the single solution u = 1 of the quadratic equation. Good work so far. Now we have to undo the substitution. Since we originally let u = sin x, now we substitute the sine of x back to get: \(u = 1\) \(\sin x = 1\)

OpenStudy (mathstudent55):

Now we need to solve the trig part of the equation. For which values of x is \(\sin x = 1\)?

OpenStudy (kydawg):

x= pi/2, 5pi/2, 9pi/2, -3pi/2, -7pi/2 and so on. It says list them all and there are unlimited.

OpenStudy (mathstudent55):